# Solvent: liquidity verification of smart contracts

### v2

@@ 1 Introduction
Formal methods provide an ideal defense against these attacks, since they enable the creation of tools to detect bugs in smart contracts before they are deployed. Indeed, smart contracts verification tools, often based on formal methods, have been mushrooming in the last few years: for Ethereum alone — largely the main smart contract platform — dozens of tools exist today [23]. Still, the actual effectiveness of these tools in countering real-world attacks is debatable: indeed, attacks to smart contracts have continued to proliferate, refining their strategies from exploits of known vulnerability patterns to sophisticated attacks to the contracts’ business logics. As a matter of fact, the vast majority of the losses due to real-world attacks are caused by logic errors in the contract code [13], which are outside the scope of most vulnerability detection tools. In particular, several real-world attacks were based on liquidity weaknesses of smart contracts, which were exploited by attackers to steal or freeze crypto-assets [3]. Liquidity expresses the ideal behaviour of contracts in terms of the exchange of crypto-assets [34, 10, 24]: users want to be guaranteed that, whenever certain states are reached, they can always perform some actions that lead to a desirable assets transfer [29]. The resulting class of properties is dubbed, in general, liquidity.
@@
The current tools that support the verification of behavioural properties of Solidity contracts cannot verify general liquidity properties. This is due both to the design choices of Solidity, their target language, and to the complex logical structure of such properties. The difficulty of verifying Solidity is caused by several glitches of high-level abstractions over the low-level target (e.g. reentrancy [5], gas costs [18], and non-native tokens [36]). Indeed, it has been observed that existing tools already face challenges (in terms of soundness and completeness) in the verification of even simpler properties [9]. We believe that effectively verifying liquidity requires first to purify Solidity from its main semantical quirks. Once we have done this, we must deal with the peculiar logical structure of liquidity properties. We start by noting that liquidity cannot be expressed in terms of safety or liveness (see also Section 0.A.2). Indeed, while safety has the form “$p$ always holds”, and liveness has the form “$p$ eventually holds”, liquidity instead has the form “for certain users $a$, there always exists a sequence of at most $m$ actions that $a$ can perform to make $p$ hold” (where $m$ is a parameter). Notice that bounded liveness cannot model liquidity either: not being able to branch over paths, it can only express properties of the form “$p$ always holds after at most $m$ steps”. Indeed, this may even be undesirable: for example, a bank client does not want to “always withdraw their deposit after at most $m$ steps” (maybe they want to keep their money deposited for $m+1$ steps); rather, they want to know whether they “can always withdraw their deposit, by performing at most $m$ actions”. To deal with liquidity, verification techniques that go beyond safety and (bounded) liveness are therefore necessary.
Contributions
@@
We propose Solvent  [^1] https://github.com/AngeloFerrando/Solvent, a tool that verifies liquidity properties of smart contracts. Solvent takes as input a contract, written in a purified version of Solidity, and a set of user-defined liquidity properties. The tool translates them into SMT constraints [8], reducing the verification problem to a symbolic model checking one. Then, techniques such as bounded model checking and predicate abstraction are employed, relying on Z3 [14] and cvc5 [6] as back-end SMT solvers. Experiments on a benchmark of real-world smart contracts show that Solvent can efficiently verify relevant liquidity properties of their behaviour. These properties are currently out of the scope of industrial verification tools operating on the full Solidity, like e.g. SolCMC [4] and Certora [21], as well as academic tools [22, 19, 28, 25, 31, 35] (due to space constraints, we discuss such tools in Section 0.A.3). Solvent provides developers with useful feedback, by detecting logical errors that would otherwise remain unnoticed. In particular, when Solvent detects a property violation, it produces a concrete execution trace that leads the contract to a state from which the desired asset exchange is unrealisable.
@@
  • a thorough evaluation of the tool effectiveness and performance on a benchmark of real-world smart contracts, which we extend with relevant liquidity properties (available on the tool github repository);
  • a concrete demonstration of the tool applicability, showing subtle bugs in existing smart contracts that cause crypo-assets to get frozen forever.
@@ 2 Verifying liquidity properties with Solvent
require (target_reached && block.number>end_donate);
@@
To translate donor_wd into an SMT constraint, Solvent considers its negation, and introduces a new existentially quantified variable for xa, and new universally quantified variables for all transaction variables in the sequence tx and for all next-state variables. Then, it reduces to checking if there exists an xa for which, if the antecedent holds, for all transaction variables and next-state variables, either the transactions invalidly tie current and next-state, or the required consequent does not hold. If this formula is unsatisfiable, then the property holds; otherwise, a counterexample is produced. E.g., for donors_wd, the counterexample is the following sequence of transitions. [^2] Transition 3 represents a call to an adversarial contract, with a method invoking selfdestruct on the crowdfund contract. We note that selfdestruct is still active [16]. This can be easily translated into a concrete PoC, leading the contract to a state where the property cannot be satisfied.
@@ 3 Evaluation
We test our tool over a common benchmark for Solidity verification [2], which includes a representative set of real-world contracts and properties. Since this benchmark is focused on current verification tools for Solidity, which do not deal with general liquidity properties, we extend it with relevant properties of this kind for each contract (see the github page). Overall, we end up with 107 verification tasks, which we manually check for the ground truth.
@@
We run Solvent on each verification task on a 3GHz 64-bit Intel Xeon Gold 6136 CPU and a GNU/Linux OS (x86_64-linux) with 64 GB of RAM, with either cvc5 (v. 1.1.3-dev.152.701cd63ef) or Z3 (v. 4.13.0) as a back-end. The run-time limit for each verification task is 1000s of CPU time. A subset of the results are shown in Table 1 (see github for the full results). We mark each property as: “✗(N)”, if the solver finds a trace that violates the property (with $N$ being the length of the shortest trace leading to a violation); “✓()”, if it proves that the property holds in all possible states; “✓(N)”, if it proves that the property holds for every trace of length at most $N$; and “?” if it timeouts.
@@ Results.
First, we note that both solvers never return an inconsistent answer. For all non-liquid properties, except one, at least one of the solvers is able to find a counterexample. When a counterexample is found, the result is returned quite quickly, and the trace is quite short. For liquid properties, the solvers are able to prove the property only for some instances. Still, in most cases, they manage to verify the property up-to traces of significant length. Two contracts (“Payment splitter” and “Vesting wallet”) are significantly tough for both solvers. This is not surprising though, as they both present non-linear behaviour, thus requiring to solve SMT formulas in the theory of Nonlinear Integer Arithmetic, which is undecidable and notoriously hard to deal with in practice for SMT solvers. [^3] Strategies to overcome the obstacles posed by NIA have been recently discussed, for the specific case of formulas coming from the verification of smart contracts, in [20].
@@ Discussion.
Auction seller_wd   ✓(10) ✓(12)
@@
no_frozen_funds ✗(3) 2.62 ✗(3) 2.59
Bank deposit_not_revert ✓() 4.88 ✓() 23.34
@@
no_frozen_funds ✗(2) 1.90 ✗(2) 2.01
Crowdfund (bug) owner_wd ✓() 4.65 ✓() 22.97
donor_wd ✗(3) 3.10 ✗(3) 42.80
no_frozen_funds ✗(1) 0.99 ✗(1) 0.90
@@
anyone_wd ✗(2) 1.77 ✗(2) 1.96
no_frozen_funds ✗(3) 2.92 ✗(3) 2.75
@@
no_frozen_funds ✓() 4.96 ✓() 29.42
Lottery (0.A.4) one_player_win ✓(11) ✓(15)
@@
wd_fin_owner ✗(1) 1.34 ✗(1) 1.18
Vesting wallet owner_wd_expired ✓() 6.46 ? T/O
@@
owner_wd_uncond ? T/O ✗(1) 1.14
owner_wd_beforestart ? T/O ✗(1) 0.96
owner_wd_empty ? T/O ✗(1) 0.96
@@ 4 Conclusions and Future work
Solvent has already proven useful to spot general liquidity vulnerabilities, which are beyond the reach of current verification tools. Still, there is space for improvements. First, Solvent uses off-the-shelf SMT solvers, relying on bounded model checking to find counterexamples, and on predicate abstraction to prove that the property holds. An alternative approach is to leverage advanced techniques (e.g., abstraction-refinement, k-induction, etc.) used by modern infinite-state symbolic model checkers. Further work involves extending the Solvent fragment to narrow the gap with the actual Solidity. This could amount to extend the supported Solidity fragment to include e.g. contract-to-contract calls (see Appendix A for further extensions), and to automatically transform the transaction sequences outputted by Solvent into an executable PoC in the HardHat tool. The formalization of liquidity properties in a temporal logic is left as future work.
Acknowledgments
@@ Appendix 0.A Supplementary material
  • in Section 0.A.2 we discuss why liquidity properties cannot be expressed as safety or liveness properties
  • in Section 0.A.3 we discuss related work on bug detection and formal verification tools for smart contracts. In particular, we provide a thoroughly discussion of the differences between Solvent and the two main verification tools for Solidity, i.e. SolCMC and Certora, highlighting why verification of liquidity properties is currently out of their scope. We also discuss other verification tools, such as VeriSolid [25] and SmartPulse [31], that can verify liveness properties but not liquidity ones.
  • in Section 0.A.4 we evaluate Solvent on a more complex use case, i.e. a fair 2-players lottery. Implementing this use case is quite error-prone, since the contract must implement a commit-reveal protocol that prescribes punishments whenever a player behaves dishonestly, e.g. by refusing to perform some required action. The contract must ensure that, even in these cases, an honest player has at least the same payoff that she would have by interacting with another honest player. Proving relevant liquidity properties of this contract seems far beyond the capabilities of existing verification tools. To overcome these limitations, we provide Solvent with an extended syntax to express hashed timelock protocols, similarly to what is done in other smart contract languages (e.g., in Tezos).
@@ 0.A.1 Background on Solidity, and relation to Solvent’s fragment
  • In Solidity, transfers of crypto-currency from a contract to another account are encoded as contract-to-contract calls, while in Solvent we assume that the transferred amount actually arrives at the destination address, i.e. that this address is either an EOA or a contract account that does not perform a further call. This assumption simplifies stating properties about the funds transferred from a contract (see Section 0.A.3), and is used in other verification tools such as VerX [28], where it is called effective external callback freedom assumption.
@@
  • Solvent features a special syntax to express hashed timelock protocols, where a user first store the hash of a chosen secret in the contract, and then reveal the secret, making the contract check that the hash of the revealed secret corresponds to the stored hash. This is a common pattern, used e.g. in gambling games and blind auctions, but where existing verification tools for Solidity fail to give the expected results (see Section 0.A.3). We exemplify this feature to design a 2-players lottery in Section 0.A.4, showing that Solvent manages to provide developers with a useful feedback.
0.A.2 Liquidity vs. Safety and Liveness
In this section, we support our claim that liquidity properties cannot be expressed in terms of safety or liveness. First, we provide an intuitive motivating example; then, we discuss the structural differences between liquidity and safety and liveness properties in terms of LTL and CTL.
Consider a user opening a bank account. The user wants to be guaranteed that, if they deposit a certain amount of money in the bank, then they will always be able to withdraw their money by performing a reasonable amount of actions, not that they always withdraw their money, nor that they will eventually withdraw their money. Requiring liveness but not liquidity would allow the bank to give a guarantee such as “any client will withdraw their money after the bank has been visited overall 1 million times after the deposit”, making the liveness property “eventually I will withdraw” true (assuming the fairness condition that visits to the bank happen infinitely often). But it does not make true the more reasonable and desirable liquidity property “I’m always able to perform at most 3 actions that make me withdraw”. Note that bounded liveness would not be an appropriate specification either. Indeed, the user does not want to “always withdraw after 3 actions are made” (the 3 actions could be “deposit more money”, or they could be made by other users). Branching is necessary, as we need to existentially quantify over the paths.
In terms of LTL, we have that safety properties have the form “${G}p$”, and liveness properties have the form “${GF}p$”. Liquidity properties, however, cannot be expressed in LTL, as we need to existentially quantify over the paths. In CTL, the closest formulation would be “${AG\ EX\ p\lor AG\ EX\ EX\ p\lor\dots\lor AG\ (EX)^{m}\ p}$” (with $m$ fixed). This formulation, however, still does not capture the strategic aspect of liquidity, as we require that the sequence that makes $p$ hold consists of actions made by a certain user, and therefore we need to impose conditions over transition variables.
0.A.3 Related work
The Ethereum ecosystem includes several bug detection tools that can spot specific forms of liquidity vulnerabilities in smart contracts. In that setting, these vulnerabilities are often referred to as “Locked Ether”, which are roughly described as the absence of a mechanism to withdraw Ether from the contract. The tools capable of detecting Locked Ether bugs include Slither [17], SmartCheck [32], Maian [27], Securify2 [34], ConFuzzius [33] and sFuzz [26]. Each of these tools has its own internal encoding of the Locked Ether property, making it difficult to compare them [30]. There are two main differences between these tools and ours. First, while these tools have an hard-coded Locked Ether property, Solvent can verify custom liquidity properties, defined by our specification language. Second, bug detection tools only focus on property violations, while Solvent (a formal verification tool) in some cases can successfully verify that a liquidity property is satisfied.
Current formal verification tools for Solidity do not support general liquidity properties. SolCMC, the prover shipped with the Solidity compiler, and other verification tools such as Zeus [22], solc-verify [19], VerX [28] and SmartACE [35], only support safety properties, and therefore they cannot express any of the liquidity properties considered in this paper.
Other tools such as Certora [21], VeriSolid [25], and SmartPulse [31], can also verify liveness properties. In the rest of this section we discuss more concretly why Certora, SmartPulse and VeriSolid cannot express general liquidity properties.

### v3

@@ 1 Introduction
Formal methods provide an ideal defense against these attacks, since they enable the creation of tools to detect bugs in smart contracts before they are deployed. Indeed, smart contracts verification tools, often based on formal methods, have been mushrooming in the last few years: for Ethereum alone — largely the main smart contract platform — dozens of tools exist today [26]. Still, the actual effectiveness of these tools in countering real-world attacks is debatable: indeed, attacks to smart contracts have continued to proliferate, refining their strategies from exploits of known vulnerability patterns to sophisticated attacks to the contracts’ business logics. As a matter of fact, the vast majority of the losses due to real-world attacks are caused by logic errors in the contract code [15], which are outside the scope of most vulnerability detection tools. In particular, several real-world attacks were based on liquidity weaknesses of smart contracts, which were exploited by attackers to steal or freeze crypto-assets [3]. Liquidity (also called solvency [25] or enabledness [32]) expresses the ideal behaviour of contracts in terms of the exchange of crypto-assets [37, 12, 27]: users want to be guaranteed that, whenever certain states are reached, they can always perform some actions that lead to a desirable assets transfer. There are two key points in this notion: the user wants to be able to constrain who can perform these actions, and how many actions are needed [32].
@@
The current tools that support the verification of Solidity cannot verify general liquidity properties. This is due both to the design choices of Solidity, their target language, and to the complex logical structure of such properties. The difficulty of verifying Solidity is caused by several glitches of high-level abstractions over the low-level target (e.g. reentrancy [5], gas costs [20], and non-native tokens [39]). Indeed, it has been observed that existing tools already face challenges (in terms of soundness and completeness) in the verification of even simpler properties [11]. We believe that effectively verifying liquidity requires first to purify Solidity from its main semantical quirks. Once we have done this, we must deal with the peculiar logical structure of liquidity properties. Indeed, liquidity cannot be expressed in terms of safety or liveness, and bounded liveness cannot model liquidity either (see  Section 2). To deal with liquidity, verification techniques that go beyond safety and (bounded) liveness are therefore necessary.
Contributions.
@@
We propose Solvent, a tool that verifies liquidity properties of smart contracts. Solvent takes as input a contract, written in a purified version of Solidity, and a set of user-defined liquidity properties. The tool translates them into SMT constraints [9], reducing the verification problem to an SMT-based symbolic model checking one [8]. Then, techniques such as bounded model checking and predicate abstraction are employed, relying on Z3 [16] and cvc5 [6] as back-end SMT solvers. Experiments on a benchmark of real-world smart contracts show that Solvent can efficiently verify relevant liquidity properties of their behaviour. These properties are currently out of the scope of industrial verification tools operating on the full Solidity, like e.g. SolCMC [4] and Certora [23], as well as academic tools [24, 21, 31, 28, 34, 38] (we discuss such tools in Section 5). Solvent provides developers with useful feedback, by detecting logical errors that would otherwise remain unnoticed. In particular, when Solvent detects a property violation, it produces a concrete execution trace that leads the contract to a state from which the desired asset exchange is unrealisable.
@@
  • a thorough evaluation of the tool effectiveness and performance on a benchmark of real-world smart contracts, which we extend with relevant liquidity properties (available on the tool github repository [10]);
  • a concrete demonstration of the tool applicability, showing subtle bugs in existing smart contracts that cause crypto-assets to get frozen forever.
2 On liquidity
In this section, we clarify our notion of liquidity, and we support our claim that it cannot be expressed in terms of safety or (bounded) liveness. Recall that safety has the form “$p$ always holds”, liveness has the form “$p$ eventually holds”, and bounded liveness “in at most $m$ steps $p$ holds”. Liquidity, instead, has the form “for certain users $a$, there always exists a sequence of at most $m$ actions that $a$ can perform to make $p$ hold” (where $m$ is a parameter).
To illustrate the differences among these notions, consider a user opening a bank account. The user wants to be guaranteed that, if they deposit money in the bank, then they will always be able to withdraw their money by performing a reasonable amount of actions. They do not want to always withdraw their money (safety), or to just eventually withdraw their money (liveness). For example, requiring liveness but not liquidity would allow the bank to give a guarantee such as “any client will withdraw their money after the bank has been visited overall 1 million times after the deposit”, making the liveness property “eventually I will withdraw” true (assuming the fairness condition that visits to the bank happen infinitely often). However, the more desirable liquidity property “I am always able to perform at most 3 actions that make me withdraw” would not hold. Bounded liveness would not be an appropriate requirement either. Indeed, assume that the bank guarantees that “the user always receives their money before 3 actions have been made”. This would be undesirable, since it says that any sequence of 3 actions, performed by any user, would trigger the withdrawal.
In LTL, we have that safety properties have the form “${G}p$”, and liveness properties have the form “${GF}p$”. Liquidity properties, however, cannot be expressed in LTL, as we need to existentially quantify over the paths (“there exists a path where the user performs at most $m$ actions to make $p$ hold”). In CTL, the closest formulation would be “${AG\ EX\ p\lor AG\ EX\ EX\ p\lor\dots\lor AG\ (EX)^{m}\ p}$”. This, however, still does not capture the strategic aspect of liquidity, which requires that the sequence that makes $p$ hold consists of actions made by certain users, and therefore we need to impose conditions over transition variables.
@@ 3 Verifying liquidity properties with Solvent
require (block.number>end_donate && target_reached);
@@
To translate donor_wd into an SMT constraint, Solvent considers its negation, and introduces a new existentially quantified variable for xa, and new universally quantified variables for all transaction variables in the sequence tx and for all next-state variables. Then, it reduces to checking whether there exists an xa for which, if the antecedent holds, for all transaction variables and next-state variables, either the transactions invalidly tie current and next-state, or the required consequent does not hold. If this formula is unsatisfiable, then the property holds; otherwise, a counterexample is given. E.g., for donors_wd, the counterexample is the following sequence of transitions:
@@
Here, the last transition represents a call to an adversarial contract, with a method invoking selfdestruct on Crowdfund. [^1] Note that selfdestruct is still active [18]. This can be easily translated into a concrete Proof-of-Concept (PoC), leading the contract to a state where donors_wd cannot be satisfied.
@@
Note that the fixed contract still has a liquidity vulnerability: if the target is not reached, donors can redeem their donations, but any extra funds possibly existing at contract creation, or received through block rewards or selfdestruct actions, will be frozen in the contract. To fix this vulnerability, we first need to quantify these extra funds: we do this by adding a variable tot_donations that we update in donate and in wdDonor, so that the extra budget is now given by balance-tot_donations. We then add a method that allows anyone to transfer any extra budget to the owner after the deadline. Solvent verifies that the revised contract still enjoys donor_wd, and transfers any extra budget to the owner, i.e.:
setfontsize89
property no_frozen_funds {
Forall xa [ balance>tot_donations && block.number>end_donate
-> Exists tx [1, xa]
[ (<tx>balance[owner] >= balance[owner]
+ (balance - tot_donations)) ] ] }
[…]
@@ 4 Evaluation
We test our tool over a common benchmark for Solidity verification [2], which includes a representative set of real-world contracts and properties. Since this benchmark is focussed on current verification tools for Solidity, which do not deal with general liquidity properties, we extend it with relevant properties of this kind for each contract (see the github page). Overall, we end up with 107 verification tasks, which we manually check for the ground truth.
@@
We run Solvent on each verification task on a 3GHz 64-bit Intel Xeon Gold 6136 CPU and a GNU/Linux OS (x86_64-linux) with 64 GB of RAM, with either cvc5 (v. 1.1.3-dev.152.701cd63ef) or Z3 (v. 4.13.0) as a back-end. The run-time limit for each verification task is 400s of CPU time. A subset of the results are shown in Table 1 (see github for the full results). We mark each property as: “✗(N)”, if the solver finds a trace that violates the property (with $N$ being the length of the shortest trace leading to a violation); “✓()”, if it proves that the property holds in all possible states; “✓(N)”, if it proves that the property holds for every trace of length at most $N$; and “?” if it timeouts.
@@ Results.
First, we note that both solvers never return an inconsistent answer. For all non-liquid properties, except two, at least one of the solvers is able to find a counterexample. When a counterexample is found, the result is returned quite quickly, and the trace is quite short. For liquid properties, the solvers are able to prove the property only for some instances. Still, in most cases, they manage to verify the property up-to traces of significant length. Two contracts (“Payment splitter” and “Vesting wallet”) are significantly tough for both solvers. This is not surprising though, as they both present non-linear behaviour, thus requiring to solve SMT formulas in the theory of Nonlinear Integer Arithmetic, which is undecidable and notoriously hard to deal with in practice for SMT solvers. [^2] Strategies to overcome the obstacles posed by NIA have been recently discussed, for the specific case of formulas coming from the verification of smart contracts, in [22].
@@ Discussion.
Auction no_frozen_funds ✗(3) 1.10 ✗(3) 1.15
seller_wd   ✓(10) ✓(9)
@@
Bank deposit_not_revert ✓() 2.04 ✓() 19.79
@@
no_frozen_funds ✗(2) 0.70 ✗(2) 0.74
Crowdfund (bug) owner_wd ✓() 2.31 ✓() 26.64
donor_wd ✗(3) 1.38 ✗(3) 7.78
  Crowdfund (fix2) owner_wd ✓() 2.34 ✓() 29.34
donor_wd ✓(7) ✓(6)
no_frozen_funds ✓(9) ✓(8)
@@
anyone_wd ✗(2) 0.70 ✗(2) 0.76
no_frozen_funds ✗(3) 1.17 ✗(3) 1.23
@@
no_frozen_funds ✓() 2.37 ✓() 54.84
Lottery (0.A.3) one_player_win ✓(13) ✓(9)
@@
wd_fin_owner ✗(1) 0.45 ✗(1) 0.44
Vesting wallet owner_wd_expired ✓() 2.16 ? T/O
@@
owner_wd_uncond ? T/O ✗(1) 0.36
owner_wd_beforestart ? T/O ✗(1) 0.35
owner_wd_empty ? T/O ✗(1) 0.35
@@
5 Related work
The Ethereum ecosystem includes several bug detection tools that can spot specific forms of liquidity vulnerabilities in smart contracts. In that setting, these vulnerabilities are often referred to as “Locked Ether”, which are roughly described as the absence of a mechanism to withdraw Ether from the contract. The tools capable of detecting Locked Ether bugs include Slither [19], SmartCheck [35], Maian [30], Securify2 [37], ConFuzzius [36] and sFuzz [29]. Each of these tools has its own internal encoding of the Locked Ether property, making it difficult to compare them [33]. There are two main differences between these tools and ours. First, while these tools can only spot that a predefined hard-coded Locked Ether property is violated, Solvent can verify custom liquidity properties, defined by our specification language. In particular, Solvent can verify properties that express who can withdraw how much, and under which conditions; not just that Ether does not get frozen in the contract. Second, bug detection tools only focus on property violations, while Solvent (a formal verification tool) is capable of verifying that a liquidity property is satisfied (always, or up to a certain bound on the number of transactions).
Other state-of-the-art formal verification tools for Solidity do not support general liquidity properties. SolCMC, the prover shipped with the Solidity compiler, as well as other verification tools such as Zeus [24], solc-verify [21], VerX [31] and SmartACE [38], only support safety properties, while other tools, such as Certora [23], VeriSolid [28], and SmartPulse [34], can also verify liveness properties. However, as discussed in Section 2, this is not enough for liquidity properties (see Appendix A for a more in-depth discussion).
@@ 6 Conclusions and Future work
Solvent has already proven useful to spot general liquidity vulnerabilities, which are beyond the reach of current tools. Still, there is space for improvements.
First, Solvent uses off-the-shelf SMT solvers, relying on bounded model checking to find counterexamples, and on predicate abstraction to prove that the property holds. An alternative approach that we plan to investigate is to leverage advanced techniques (e.g., abstraction-refinement, k-induction, etc.) used by modern infinite-state symbolic model checkers.
Second, we plan to extend our Solidity fragment to narrow the gap with the actual Solidity (see Appendix A for a discussion of the main differences). In particular, we note that concretising our transfer into contract-to-contract calls, as in Solidity, would require to take failures and reentrancy into account. E.g., consider a contract with a method:
setfontsize89
foo() { owner.transfer(1); msg.sender.transfer(1); }
[…]
and a property requiring that, whenever the contract balance is at least 2, any user can increase their balance by 1. In our Solidity fragment, this property is true, since both transfers will succeed under the given hypotheses. This is not the case in the concrete Solidity: e.g., owner could be a contract address, whose fallback method could fail under certain conditions. In this case, the second transfer in foo would not happen, and so the property would be violated. Since considering each transfer as potentially failing would make most contracts illiquid, a possible approach would be to allow queries to specify which transfers or addresses to be considered trusted.
Other future work include the automatic transformation of the counterexamples given by Solvent into an executable PoC (e.g., in the HardHat tool), and the formalisation of liquidity in a suitable logic.
Acknowledgments.
@@ Appendix 0.A Supplementary material
  • in Section 0.A.2 we discuss related work on bug detection and formal verification tools for smart contracts. In particular, we provide a thoroughly discussion of the differences between Solvent and the two main verification tools for Solidity, i.e. SolCMC and Certora, highlighting why verification of liquidity properties is currently out of their scope. We also discuss other verification tools, such as VeriSolid [28] and SmartPulse [34], that can verify liveness properties but not liquidity ones.
  • in Section 0.A.3 we evaluate Solvent on a more complex use case, i.e. a fair 2-players lottery. Implementing this use case is quite error-prone, since the contract must implement a commit-reveal protocol that prescribes punishments whenever a player behaves dishonestly, e.g. by refusing to perform some required action. The contract must ensure that, even in these cases, an honest player has at least the same payoff that she would have by interacting with another honest player. Proving relevant liquidity properties of this contract seems far beyond the capabilities of existing verification tools. To overcome these limitations, we provide Solvent with an extended syntax to express hashed timelock protocols, similarly to what is done in other smart contract languages (e.g., in Tezos).
@@ 0.A.1 Background on Solidity, and relation to Solvent’s fragment
  • In Solidity, transfers of crypto-currency from a contract to another account are encoded as contract-to-contract calls, while in Solvent we assume that the transferred amount actually arrives at the destination address, i.e. that this address is either an EOA or a contract account that does not perform a further call. This assumption simplifies stating properties about the funds transferred from a contract (see Section 0.A.2), and is used in other verification tools such as VerX [31], where it is called effective external callback freedom assumption.
@@
  • Solvent features a special syntax to express hashed timelock protocols, where a user first store the hash of a chosen secret in the contract, and then reveal the secret, making the contract check that the hash of the revealed secret corresponds to the stored hash. This is a common pattern, used e.g. in gambling games and blind auctions, but where existing verification tools for Solidity fail to give the expected results (see Section 0.A.2). We exemplify this feature to design a 2-players lottery in Section 0.A.3, showing that Solvent manages to provide developers with a useful feedback.
0.A.2 Verification of liquidity in other tools
We now discuss more concretely why Certora, SmartPulse and VeriSolid cannot express general liquidity properties.